Solve the problem.
Use synthetic division to divide f(x) = x3 + 1x2 - 26x + 24 by x + 6. Use the result to find all zeros of f.

Answer by lwsshak3(6761)

Solve the problem.
Use synthetic division to divide f(x) = x3 + 1x2 - 26x + 24 by x + 6. Use the result to find all zeros of f.
**
-6......1.......1.......-26........24
.................-6.........30......-24
..........1.....-5.........4...........0 (-6 is a zero)
f(x)=(x+6)(x^2-5x+4)
f(x)=(x+6)(x-4)(x-1)
zeros are: -6, 1 and 4